# Branch, or Layer? Zeroth-Order Optimization for Continual Learning of Vision-Language Models

### v1

@@ 3.4 A New Enhancement from Vision Branch Discrepancy
Table 3: Effect of vision branch discrepancy on performance. Dual w/ Hop-odd/even indicates the results of adopting Hop-odd and Hop-even layer-wise ZO in dual branch.
Method CIFAR-100 Inc10 Tiny-ImageNet Inc20 Image-R Inc20
@@
Vision Discrepancy 79.87 87.25 52.46 61.23 65.80 71.82
Gradient Regularization and Vision Branch Perturbation Control. To validate this hypothesis, we propose a ZO optimization strategy that incorporates gradient regularization during the estimation process, with explicit constraints on perturbations applied to the vision branch to mitigate instability. Specifically, we first introduce a signed gradient transformation to unify the gradient values which can be expressed as:
$\tilde{g}=\operatorname{sign}(g),\quad\text{where}\quad\operatorname{sign}(g_{i})=\begin{cases}+1,&\text{if }g_{i}>0\\ 0,&\text{if }g_{i}=0\\ -1,&\text{if }g_{i}<0\end{cases}\quad\forall i\in\{1,\dots,d\},$ (3)
where $g$ denotes the original gradient, and $\tilde{g}$ is the transformed signed gradient. This transformation retains only the direction information of the gradient and discarding the amplitude information. Furthermore, we implement modality-specific perturbation factors for ZO optimization, assigning a deliberately lower value to the vision branch compared to the language branch. This strategy theoretically reduces the instability of ZO optimization and suppresses gradient fluctuations in the vision modality. To verify this strategy, we first record the gradient variance on CIFAR-100, and plot the comparison shown in Figure 7. It can be observed that our strategy significant reduce the vision branch gradient fluctuations expressed in Figure 7(b) and 7(d) (shown in red line), regardless of Hop-odd or Hop-even. We further validate this conjecture through quantitative analysis shown in Table 3, it can be find that stabilizing ZO gradient estimation and reducing vision branch perturbations yields consistent performance gains across CIFAR-100, Tiny-Imagenet and Image-R datasets. This method focus on the instability limits of ZO in VLCL and maintains the balance of the optimization process across different modality branches, paving the way to further refines the application strategy of ZO optimization in VLCL.
@@
Figure 7: Analyzing the effect of vision discrepancy on gradient variance in Hop-odd vs. Hop-even. Blue shows original results, while red shows the positive impact attributed to vision discrepancy.
4 Ablation, Empirical Demonstration and Beyond
Exploring diverse ZO strategies. Our ablation start to explore other ZO strategies of hybrid ZO-FO collaboration in VLCL. As shown in Table 4, applying ZO to dual branches significantly degrades performance compared to the baseline, confirming the instability caused by excessive gradient oscillations. In contrast, single-branch ZO optimization mitigates this issue, with the significant improvement in Last. and Avg.. We then explore the impact of different ZO optimization on VLCL performance. As we default to the conservative ZO strategy in the main analysis, we further conduct more aggressive ZO variants. Specifically, we examine a naive ZO approach—referred to as ZO* which performs a single gradient estimation and directly updates the parameters without any loss-based validation. On top of this, we introduce a variant named Sign, which incorporates a signed gradient transformation to regulate the magnitude of the estimated gradients. This design aims to mitigate the instability caused by ZO estimates while preserving directional information.
Table 4: Behavioral consistency across diverse ZO strategies. * indicates a naive ZO strategy. And Sign indicates a ZO strategy using the gradient transformation method.
Method Strategy CIFAR-100 Inc10
@@ 4 Ablation, Empirical Demonstration and Beyond
It can be observed that the aggressive ZO* optimization degrades performance, as it relies on a single gradient estimation without validating the update direction, making it more sensitive to the fluctuations in the loss landscape compared with conservative ZO. Notably, adopting Sign-based gradient estimation enhances performance across all configurations, suggesting that enforcing gradient amplitude consistency in ZO perturbations helps stabilize the optimization trajectory and improves convergence. This observation aligns with our proposed ZO optimization strategy in Section 3.4, which advocates for gradient magnitude control as a means to suppress fluctuation updates and maintain optimization balance across modalities.
(a) Last. of CIFAR-100.
(b) Avg. of CIFAR-100.
(c) Last. of Image-R.
(d) Avg. of Image-R.
@@
Significance analysis. We then explore the average incremental performance among 5 runs on CIFAR-100 and Image-R, with 3 different ZO configurations (dual branch, vision branch and language branch), recording the accuracy of Last. and Avg. shown in Figure 8. We observe that ZO optimization yields smaller performance variance on the CIFAR-100 dataset than on Image-R across different configurations, regardless of Last and Avg. One possible explanation is that Image-R derived from the larger-scale ImageNet, contains more fine-grained and semantically diverse categories than CIFAR-100. This increased label complexity leads to a more rugged and high-dimensional loss landscape,
Table 5: Comparison of GPU memory usage across branches(CLIP).
Method Memory $\downarrow$
MoE4Adapter $\sim$19.96GB
      Dual w/ ZO $\sim$2.17GB
      Vision w/ ZO $\sim$6.93GB
      Language w/ ZO $\sim$12.39GB
MoE4Adapter† $\sim$15.11GB
      Dual w/ ZO $\sim$1.73GB
      Vision w/ ZO $\sim$5.71GB
      Language w/ ZO $\sim$11.09GB
@@
which makes ZO gradient estimation on stochastic perturbations harder to stabilize. Additionally, from the perspective of different ZO configurations, applying ZO optimization to the language branch consistently yields the better performance. We hypothesize that this is because the language branch typically operates on lower-dimensional tensors, making it less susceptible to the instability introduced by random perturbations in ZO gradient estimation. In contrast, the higher-dimensional vision branch benefits more from FO gradients, which provide precise gradient optimization direction. The combination of FO-dominated updates in the vision branch and ZO exploration in the language branch may facilitate better escape from local minima and lead to more robust convergence.
@@
Quantifying memory efficiency. We record the memory usage across different ZO settings, as shown in Table 5. Compared to applying FO optimization in the dual branch MoE4Adapter, using ZO optimization leads to a 89.1% reduction in memory consumption, due to the elimination of gradient backpropagation and storage overhead. When applying ZO optimization to a single branch, memory consumption is reduced by 65.3% on the visual branch and 37.9% on the language branch, respectively. The greater memory reduction on the visual branch is attributed to its higher input dimensionality, as it processes image data. Applying ZO in this context helps to alleviate memory pressure by avoiding storage of high-dimensional gradients. For the results in LoRA architecture, it inherently requires fewer trainable units to perform parameters update, resulting in overall lower memory consumption. Overall, these findings highlight ZO’s inherent advantage in significantly reducing memory consumption, especially in high-dimensional settings, making it a highly practical and efficient optimization alternative for resource-constrained VLCL scenarios where memory efficiency is critical.
5 Related Work
@@
Continual learning of VLM. Recent advances in CL have witnessed the emergence of VLMs as promising solutions to mitigate catastrophic forgetting through their generalized multimodal representations. Pioneering studies lwf-vr ; clip_cl2 demonstrate that pretrained VLMs like CLIP clip inherently possess remarkable continual learning capabilities even without fine-tuning. PROOF vlm_lwf further enhances CL robustness by integrating multimodal cues with adaptive mapping strategies. Nevertheless, conventional VLM-based approaches predominantly emphasize task-specific feature acquisition for new domains ralf , inadvertently compromising the integrity of previously learned representations and leading to progressive performance degradation. To address this limitation, recent efforts adopt PEFT techniques peft1 ; peft2 that selectively update lightweight modules while maintaining frozen backbone parameters. MoE4Adapters MoE4Adapters introduces a mixture-of-experts (MoE) architecture, enabling task-specific feature specialization without cross-task interference and achieving state-of-the-art performance. However, these methods universally rely on FO optimization, inherently restricting their capacity to explore optimal parameter trajectories during optimization.
@@
Optimization for continual learning. From an optimization view, existing CL methods predominantly focus on reconciling the stability-plasticity dilemma staplast through gradient technology unigrad . Orthogonal parameter updates grad1 ; grad2 ; grad3 ; grad4 and sharpness-aware minimization flat1 ; flat2 ; flat3 represent two mainstream directions, which respectively aim to decouple task-specific gradients and converge to flat loss minima for improved generalization. While these strategies enhance optimization stability, they often inadequately address the exploration-exploitation trade-off, as deterministic gradient descent trajectories tend to converge to suboptimal local minima with limited perturbation resilience. Emerging ZO optimizationzo1 ; zo2 ; zeroflow presents a paradigm shift by employing gradient-free stochastic perturbations to estimate descent directions. This approach offers dual advantages: 1) inherent stochasticity facilitates escape from local optima through controlled parameter space exploration, and 2) elimination of backward propagation reduces memory overhead by avoiding gradient matrix storage.
@@ 6 Conclusion
Limitation. This work currently focuses on CLIP-based architectures and image-text modalities. The generalization of ZO-FO collaboration to other VLMs (e.g., multimodal transformers with audio or video inputs) remains unexplored, particularly in scenarios where modalities exhibit heterogeneous feature distributions or temporal dependencies.
@@
In the main paper, we analyze the application of ZO optimization in VLCL, covering both branch-wise and layer-wise settings. Extensive evaluations are conducted based on the MoE4Adapter baseline, and further preliminary explorations are carried out under the LoRA-based setting. In the supplementary, we provide additional results to examine the consistency of our observations under the LoRA setting. This material is divided into two sections:
  • Section A investigates the performance of layer-wise allocation of ZO and FO optimization under the LoRA setting, building upon the performance trends and layer heterogeneity analyses presented in the main paper.
  • Section B explores the vision branch discrepancy of ZO optimization and evaluates the proposed method from the main paper within the LoRA setting.
@@ Appendix A Investigating Layer-wise ZO optimization for VLCL under LoRA Setting
We give the performance of layer-wise ZO optimization with both dual and single modality branches under LoRA setting shown in Table 6. Specifically, we choose two layer configurations: Hop-odd and Suffix (six). It can be observed that the performance trends are consistent with those reported in the main paper, demonstrating the layer-wise ZO optimization remains effective under the LoRA setting. We also show the convergence behavior of Hop-odd on Image-R dataset in Figure 9, and find that the training process remains consistently stable throughout.
Table 6: How ZO optimization affects VLCL through different layers (CLIP) under LoRA based setting (represented by †). We choose two configurations across layers from different branches: w/ Hop-odd and w/ Suffix (six).
Method CIFAR-100 Inc20 CIFAR-100 Inc10 Tiny-Imagenet Inc20 Image-R Inc20
@@
(a) Dual w/ Hop-odd †.
(b) Vision w/ Hop-odd †.
(c) Language w/ Hop-odd †.
Figure 9: Analyzing convergence behavior of VLCL in Hop-odd across Dual, Vision, Language under LoRA setting (represented by †).
Additionally, we similarly observe that interleaved layer-wise ZO manner (Hop-odd) yields superior performance compared with consecutive layers under the LoRA setting. To better understand the results, we separately record the gradient distributions of the Hop-odd and Hop-suf layer-wise strategies across the dual, vision, and language branches on the Image-R dataset. Figure 10 clearly reveals that interleaved layer-wise ZO optimization yields more stable gradient behavior during training, suggesting interleaving of ZO and FO better meet each layer’s needs for gradient exploration and stability, which aligns with the hypothesis proposed in the main paper.
(a) Dual w/ Hop-suf †.
(b) Vision w/ Hop-suf †.
(c) Language w/ Hop-suf †.
(d) Dual w/ Hop-odd †.
(e) Vision w/ Hop-odd †.
(f) Language w/ Hop-odd †.
Figure 10: How ZO optimization affects gradient variance across layers under LoRA setting (represented by †).
Appendix B Exploring ZO Vision Branch Discrepancy under LoRA Setting
In the main paper, we explore the discrepant behavior of ZO across different modality branches, which motivates our proposed improvements for applying ZO in the VLCL setting. We replace the baseline architecture with LoRA and conduct the same investigation under identical experimental settings to revisit this phenomenon. Table 7 demonstrates that applying gradient regularization and reducing the perturbation in the vision branch leads to further improvements in most experimental results, this indicates that our proposed method remains effective under the LoRA setting. We further analyze this method by comparing the gradient variance in Hop-odd vs. Hop-even on image-R dataset shown in Figure 11, it can be observed that the proposed method has a positive impact, as the gradient variance becomes more stable. These results are broadly consistent with the conclusions presented in the main paper, further demonstrating the effectiveness of the proposed method.
Table 7: Effect of vision branch discrepancy on performance under LoRA setting (represented by †). Dual w/ Hop-odd/even indicates the results of adopting Hop-odd and Hop-even layer-wise ZO in dual branch.
Method CIFAR-100 Inc10 Tiny-ImageNet Inc20 Image-R Inc20
@@ Appendix B Exploring ZO Vision Branch Discrepancy under LoRA Setting
Vision Discrepancy 79.30 86.23 53.54 61.99 65.89 71.98

### v2

@@ A New Enhancement from Vision Discrepancy
Table 3: Effect of MoZO optimization on performance. Dual w/ Hop-odd/even indicates the results of adopting Hop-odd and Hop-even layer-wise ZO in dual branch.
Method CIFAR Inc10 TinyImg Inc20 ImgR Inc20
@@
MoZO 79.87 87.25 52.46 61.23 65.80 71.82
Gradient Regularization and Vision Branch Perturbation Control. To validate this hypothesis, we propose a MoZO optimization strategy that incorporates gradient regularization during the estimation process, with explicit constraints on perturbations applied to the vision branch to mitigate instability. Specifically, we first introduce a signed gradient transformation to regularize the ZO-estimated gradient $\hat{g}_{ZO}$. This can be expressed as:
$\tilde{g}=\operatorname{sign}(\hat{g}_{ZO}),$ (3)
where $\hat{g}_{ZO}$ denotes the original ZO-estimated gradient, and $\tilde{g}$ is the transformed signed gradient. The $\operatorname{sign}(\cdot)$ function is applied element-wise, defined as:
$\operatorname{sign}(x_{i})=\begin{cases}+1.&\text{if }x_{i}>0\\ 0.&\text{if }x_{i}=0\\ -1.&\text{if }x_{i}<0\end{cases}$ (4)
This transformation retains only the direction information of the gradient, discarding the amplitude information. Furthermore, we implement modality-specific perturbation factors for ZO optimization, assigning a deliberately lower value to the vision branch ($\epsilon_{v}$) compared to the language branch ($\epsilon_{l}$). Hence, the final MoZO update rule can be expressed as:
$\theta_{m,t+1}=\begin{cases}\theta_{m,t}-\eta_{t}\cdot\tilde{g}(\theta_{m,t},\epsilon_{v}\xi_{t}),&\text{if }m=\text{vision}\\ \theta_{m,t}-\eta_{t}\cdot\tilde{g}(\theta_{m,t},\epsilon_{l}\xi_{t}),&\text{if }m=\text{language}\end{cases}$ (5)
where $\theta$ is the gradient parameter, $\eta_{t}$ is the learning rate, and $\xi_{t}$ represents a perturbation vector. To verify this strategy, we first record the gradient variance on CIFAR, and plot the comparison shown in Figure 7. It can be observed that our strategy significant reduce the vision branch gradient fluctuations expressed in Figure 7(b) and 7(d) (shown in red line), regardless of Hop-odd or Hop-even. We further validate this conjecture through quantitative analysis shown in Table 3, it can be find that stabilizing ZO gradient estimation and reducing vision branch perturbations yields consistent performance gains across CIFAR, TinyImg and ImgR datasets. This method focus on the instability limits of ZO in VLCL and maintains the balance of the optimization process across different modality branches, paving the way to further refines the application strategy of ZO optimization in VLCL.
@@
Figure 7: Analyzing the effect of MoZO optimization in Hop-odd vs. Hop-even. Blue shows original results, while red shows the positive impact of vision discrepancy.
Ablation, Demonstration and Beyond
Exploring diverse ZO strategies. Our ablation start to explore other ZO strategies of hybrid ZO-FO collaboration in VLCL. As shown in Table 4, applying ZO to dual branches significantly degrades performance compared to the baseline, confirming the instability caused by excessive gradient oscillations. In contrast, single-branch ZO optimization mitigates this issue, with the significant improvement in Last. and Avg. metrics. We then explore the impact of different ZO optimization on VLCL performance. As we default to the conservative ZO strategy in the main analysis, we further conduct more aggressive ZO variants. Specifically, we examine a naive ZO approach—referred to as ZO* which performs a single gradient estimation and directly updates the parameters without any loss-based validation. On top of this, we introduce a variant named Sign, which incorporates a signed gradient transformation to regulate the magnitude of the estimated gradients. This design aims to mitigate the instability caused by ZO estimates while preserving directional information. It can be observed that the aggressive ZO* optimization degrades performance, as it relies on a single gradient estimation without validating the update direction, making it more sensitive to the fluctuations in the loss landscape compared with conservative ZO.
Table 4: Behavioral consistency across diverse ZO strategies. * indicates a naive ZO strategy, Sign indicates a ZO strategy using the gradient transformation method.
Method Strategy CIFAR Inc10
@@
Notably, adopting Sign-based gradient estimation enhances performance across all configurations, suggesting that enforcing gradient amplitude consistency in ZO perturbations helps stabilize the optimization trajectory and improves convergence. This observation aligns with our proposed ZO optimization strategy in Section A New Enhancement from Vision Discrepancy, which advocates for gradient magnitude control as a means to suppress fluctuation updates and maintain optimization balance across modalities.
(a) Last. of CIFAR.
(b) Avg. of CIFAR.
(c) Last. of ImgR.
(d) Avg. of ImgR.
@@
Significance analysis. We then explore the average incremental performance among five runs on two dataset (CIFAR and ImgR), with three different ZO configurations (dual branch, vision branch and language branch), recording the accuracy of Last. and Avg. shown in Figure 8. We observe that ZO optimization yields smaller performance variance on the CIFAR dataset than on ImgR across different configurations, regardless of Last. and Avg. metrics. One possible explanation is that ImgR derived from the larger-scale ImageNet, contains more fine-grained and semantically diverse categories than CIFAR. This increased label complexity leads to a more rugged and high-dimensional loss landscape, which makes ZO gradient estimation on stochastic perturbations harder to stabilize. Additionally, from the perspective of different ZO configurations, applying ZO optimization to the language branch consistently yields the better performance. We hypothesize that this is because the language branch typically operates on lower-dimensional tensors, making it less susceptible to the instability introduced by random perturbations in ZO gradient estimation. In contrast, the higher-dimensional vision branch benefits more from FO gradients, which provide precise gradient optimization direction. The combination of FO-dominated updates in the vision branch and ZO exploration in the language branch may facilitate better escape from local minima and lead to more robust convergence.
Table 5: Comparison of GPU memory usage between different ZO settings across branches (CLIP).
Method MoE4Adapter MoE4Adapter†
Baseline $\sim$19.96GB $\sim$15.11GB
Dual w/ ZO $\sim$2.17GB $\downarrow$ $\sim$1.73GB $\downarrow$
Vision w/ ZO $\sim$6.93GB $\downarrow$ $\sim$5.71GB $\downarrow$
Language w/ ZO $\sim$12.39GB $\downarrow$ $\sim$11.09GB $\downarrow$
@@
Quantifying memory efficiency. We record the memory usage across different ZO settings, as shown in Table 5. Compared to applying FO optimization in the dual branch baseline, using ZO optimization leads to a 89.1% reduction in memory consumption, due to the elimination of gradient backpropagation and storage overhead. When applying ZO optimization to a single branch, memory consumption is reduced by 65.3% on the visual branch and 37.9% on the language branch, respectively. The greater memory reduction on the visual branch is attributed to its higher input dimensionality, as it processes image data. Applying ZO in this context helps to alleviate memory pressure by avoiding storage of high-dimensional gradients. For the results in LoRA architecture, it inherently requires fewer trainable units to perform parameters update, resulting in overall lower memory consumption. Overall, these findings highlight ZO’s inherent advantage in significantly reducing memory consumption, especially in high-dimensional settings, making it a highly practical and efficient optimization alternative for resource-constrained VLCL scenarios where memory efficiency is critical.
Related Work
@@
Continual learning of VLM. Recent advances in CL have witnessed the emergence of VLMs as promising solutions to mitigate catastrophic forgetting through their generalized multimodal representations. Pioneering studies (lwf-vr; clip_cl2) demonstrate that pretrained VLMs like CLIP (clip) inherently possess remarkable continual learning capabilities even without fine-tuning. PROOF (lwf-vr) further enhances CL robustness by integrating multimodal cues with adaptive mapping strategies. Nevertheless, conventional VLM-based approaches predominantly emphasize task-specific feature acquisition for new domains (ralf), inadvertently compromising the integrity of previously learned representations and leading to progressive performance degradation. To address this limitation, recent efforts adopt PEFT techniques (peft1; peft2) that selectively update lightweight modules while maintaining frozen backbone parameters. MoE4Adapters (MoE4Adapters) introduces a mixture-of-experts (MoE) architecture, enabling task-specific feature specialization without cross-task interference and achieving state-of-the-art performance. However, these methods universally rely on FO optimization, inherently restricting their capacity to explore optimal parameter trajectories during optimization.
@@
Optimization for continual learning. From an optimization view, existing CL methods predominantly focus on reconciling the stability-plasticity dilemma (staplast) through gradient technology (unigrad). Orthogonal parameter updates (grad1; grad2; grad3; grad4) and sharpness-aware minimization (flat1; flat2; flat3) represent two mainstream directions, which respectively aim to decouple task-specific gradients and converge to flat loss minima for improved generalization. While these strategies enhance optimization stability, they often inadequately address the exploration-exploitation trade-off, as deterministic gradient descent trajectories tend to converge to suboptimal local minima with limited perturbation resilience. Emerging ZO optimization(zo1; derivative) presents a paradigm shift by employing gradient-free stochastic perturbations to estimate descent directions. This approach offers dual advantages: 1) inherent stochasticity facilitates escape from local optima through controlled parameter space exploration, and 2) elimination of backward propagation reduces memory overhead by avoiding gradient matrix storage.
@@ Conclusion
Limitation. This work currently focuses on CLIP-based vision-language modalities. The generalization of ZO-FO collaboration to other VLMs (e.g., multimodal transformers with audio or video inputs) remains unexplored, particularly in scenarios where modalities exhibit heterogeneous feature distributions or temporal dependencies.
Acknowledgement
This work is supported in part by the National Natural Science Foundation of China (62192783), Young Elite Scientists Sponsorship Program by CAST (2023QNRC001), and China Scholarship Council.
@@ Acknowledgement
In the main paper, we analyze the application of ZO optimization in VLCL, covering both branch-wise and layer-wise settings. Extensive evaluations are conducted based on the MoE-based setting, and further preliminary explorations are carried out under the LoRA-based setting. In the supplementary, we provide additional results to examine the consistency of our observations under the LoRA setting. This material is divided into two sections:
  • Investigating the performance of layer-wise allocation of ZO and FO optimization under the LoRA setting, building upon the performance trends and layer heterogeneity analyses presented in the main paper.
  • Exploring the vision branch discrepancy of ZO optimization and evaluates the proposed method from the main paper within the LoRA setting.
@@ Appendix A Investigating Layer-wise ZO optimization for VLCL under LoRA Setting
We give the performance of layer-wise ZO optimization with both dual and single modality branches under LoRA setting shown in Table 6. Specifically, we choose two layer configurations: Hop-odd and Suffix (six). It can be observed that the performance trends are consistent with those reported in the main paper, demonstrating the layer-wise ZO optimization remains effective under the LoRA setting. We also show the convergence behavior of Hop-odd on ImgR dataset in Figure 9, and find that the training process remains consistently stable throughout.
Table 6: How ZO optimization affects VLCL through different layers (CLIP) under LoRA based setting. We choose two configurations across layers from different branches: w/ Hop-odd and w/ Suffix (six).
Method CIFAR Inc20 CIFAR Inc10 TinyImg Inc20 ImgR Inc20
@@
(a) Du. w/Hop-odd.
(b) Vis. w/Hop-odd.
(c) Lan. w/Hop-odd.
Figure 9: Analyzing convergence behavior of VLCL in Hop-odd across Dual (Du.), Vision (Vis.), Language (Lan.) under LoRA setting.
Additionally, we similarly observe that interleaved layer-wise ZO manner (Hop-odd) yields superior performance compared with consecutive layers under the LoRA setting. To better understand the results, we separately record the gradient distributions of the Hop-odd and Hop-suf layer-wise strategies across the dual, vision, and language branches on the ImgR dataset. Figure 10 clearly reveals that interleaved layer-wise ZO optimization yields more stable gradient behavior during training, suggesting interleaving of ZO and FO better meet each layer’s needs for gradient exploration and stability, which aligns with the hypothesis proposed in the main paper.
(a) Du. w/Hop-suf.
(b) Vis. w/Hop-suf.
(c) Lan. w/Hop-suf.
(d) Du. w/Hop-odd.
(e) Vis. w/Hop-odd.
(f) Lan. w/Hop-odd.
Figure 10: How ZO optimization affects gradient variance across layers in Dual (Du.), Vision (Vis.), Language (Lan.) under LoRA setting.
Table 7: Effect of MoZO optimization on performance under LoRA setting (represented by †). Dual w/ Hop-odd/even indicates the results of adopting Hop-odd and Hop-even layer-wise ZO in dual branch.
Method CIFAR Inc10 TinyImg Inc20 ImgR Inc20
@@
MoZO 79.30 86.23 53.54 61.99 65.89 71.98
Appendix B Exploring the effect MoZO Optimization under LoRA Setting
In the main paper, we explore the discrepant behavior of ZO across different modality branches, which motivates our proposed improvements for applying ZO in the VLCL setting. We replace the baseline architecture with LoRA and conduct the same investigation under identical experimental settings to revisit this phenomenon. Table 7 demonstrates that applying gradient regularization and reducing the perturbation in the vision branch leads to further improvements in most experimental results, this indicates that our MoZO method remains effective under the LoRA setting. We further analyze this method by comparing the gradient variance in Hop-odd vs. Hop-even on CIFAR dataset shown in Figure 11, it can be observed that the proposed method has a positive impact, as the gradient variance becomes more stable. These results are broadly consistent with the conclusions presented in the main paper, further demonstrating the effectiveness of the proposed method.